DVIPS(1) DVIPS(1) NAME dvips−convertaTeXDVIfiletoPostScript SYNOPSIS dvips[options]file[.dvi] DESCRIPTION THISMANPAGEISOBSOLETE!SeetheTexinfodocumentationinstead.Youcanreadit eitherinEmacsorwiththestandaloneinfoprogramwhichcomeswiththeGNUtexinfodistributionasprep.ai.mit.edu:pub/gnu/texinfo*.tar.gz. TheprogramdvipstakesaDVIfilefile[.dvi]producedbyTeX(orbysomeotherprocessorsuch asGFtoDVI)andconvertsittoPostScript,normallysendingtheresultdirectlytothe (laser)printer.TheDVIfilemaybespecifiedwithoutthe.dviextension.Fontsusedmayeither beresidentintheprinterordefinedasbitmapsinPKfiles,ora‘virtual’combinationofboth.If themktexpkprogramisinstalled,dvipswillautomaticallyinvokeMETAFONTtogenerate fontsthatdon’talreadyexist. Formoreinformation,seetheTexinfomanualdvips.texi,whichshouldbeinstalledsomewhereon yoursystem,hopefullyaccessiblethroughthestandardInfotree. OPTIONS -aConservememorybymakingthreepassesoverthe.dvifileinsteadoftwoandonlyload-ingthosecharactersactuallyused.Generallyonlyusefulonmachineswithaverylimited amountofmemory,likesomePCs. -APrintonlyoddpages(TeXpages,notsequencepages). -bnum Generatenumcopiesofeachpage,butduplicatingthepagebodyratherthanusingthe #numcopiesoption.Thiscanbeusefulinconjunctionwithaheaderfilesetting\bophooktodocolorseparationsorotherneattricks. -BPrintonlyevenpages(TeXpages,notsequencepages). -cnum Generatenumcopiesofeverypage.Defaultis1.(Forcollatedcopies,seethe-Coption below.) -Cnum Createnumcopies,butcollated(byreplicatingthedatainthePostScriptfile).Slower thanthe-coption,buteasieronthehands,andfasterthanresubmittingthesame PostScriptfilemultipletimes. -dnum Setthedebugflags.Thisisintendedonlyforemergenciesorforunusualfact-finding expeditions;itwillworkonlyifdvipshasbeencompiledwiththeDEBUGoption.If nonzero,printsadditionalinformationonstandardoutput.Thenumberistakenasaset ofindependentbits.Themeaningofeachbitfollows.1=specials;2=paths;4=fonts; 8=pages;16=headers;32=fontcompression;64=files;128=memory;256=Kpathsea stat(2)calls;512=Kpathseahashtablelookups;1024=Kpathseapathelementexpansion; 2048=Kpathseasearches.Totraceeverythinghavingtodowithfilesearchingandopening,use3650(2048+1024+512+64+2).Totrackallclasses,youcanuse‘-1’(output isextremelyvoluminous). -Dnum Settheresolutionindpi(dotsperinch)tonum.Thisaffectsthechoiceofbitmapfonts thatareloadedandalsothepositioningoflettersinresidentPostScriptfonts.Mustbe between10and10000.Thisaffectsboththehorizontalandverticalresolution.Ifahigh resolution(somethinggreaterthan400dpi,say)isselected,the-Zflagshouldprobably alsobeused. 1June1996 1 DVIPS(1) DVIPS(1) -enum Makesurethateachcharacterisplacedatmostthismanypixelsfromits‘true’resolution-independentpositiononthepage.Thedefaultvalueofthisparameterisresolution dependent.Allowingindividualcharactersto‘drift’fromtheircorrectlyroundedpositionsbyafewpixels,whileregainingthetruepositionatthebeginningofeachnewword, improvesthespacingoflettersinwords. -EmakesdvipsattempttogenerateanEPSFfilewithatightboundingbox.Thisonly worksonone-pagefiles,anditonlylooksatmarksmadebycharactersandrules,notby anyincludedgraphics.Inaddition,itgetstheglyphmetricsfromthetfmfile,socharactersthatlieoutsidetheirenclosingtfmboxmayconfuseit.Inaddition,thebounding boxmightbeabittoolooseifthecharacterglyphhassignificantleftorrightsidebearings.Nonetheless,thisoptionworkswellforcreatingsmallEPSFfilesforequationsor tablesorthelike.(Note,ofcourse,thatdvipsoutputisresolutiondependentandthus doesnotmakeverygoodEPSFfiles,especiallyiftheimagesaretobescaled;usethese EPSFfileswithagreatdealofcare.) -fRunasafilter.Readthe.dvifilefromstandardinputandwritethePostScripttostan-dardoutput.Thestandardinputmustbeseekable,soitcannotbeapipe.Ifyoumust useapipe,writeashellscriptthatcopiesthepipeoutputtoatemporaryfileandthen pointsdvipsatthisfile.Thisoptionalsodisablestheautomaticreadingofthe PRINTERenvironmentvariable,andturnsofftheautomaticsendingofcontrolDifit wasturnedonwiththe-Foptionorintheconfigurationfile;use-Fafterthisoptionif youwantboth. -FCausesControl-D(ASCIIcode4)tobeappendedastheverylastcharacterofthe PostScriptfile.Thisisusefulwhendvipsisdrivingtheprinterdirectlyinsteadofworkingthroughaspooler,asiscommononextremelysmallsystems.NOTE!DONOTUSE THISOPTION! -hname Prependfilenameasanadditionalheaderfile.(However,ifthenameissimply‘-’sup-pressallheaderfilesfromtheoutput.)ThisheaderfilegetsaddedtothePostScriptuser-dict. -iMakeeachsectionbeaseparatefile.Undercertaincircumstances,dvipswillsplitthe documentupinto‘sections’tobeprocessedindependently;thisismostoftendonefor memoryreasons.Usingthisoptiontellsdvipstoplaceeachsectionintoaseparatefile; thenewfilenamesarecreatedreplacingthesuffixofthesuppliedoutputfilenamebya three-digitsequencenumber.Thisoptionismostoftenusedinconjunctionwiththe-S optionwhichsetsthemaximumsectionlengthinpages.Forinstance,somephototypesetterscannotprintmorethantenorsoconsecutivepagesbeforerunningoutofsteam; theseoptionscanbeusedtoautomaticallysplitabookintoten-pagesections,eachtoits ownfile. -kPrintcropmarks.Thisoptionincreasesthepapersize(whichshouldbespecified,either withapapersizespecialorwiththe-Toption)byahalfinchineachdimension.It translateseachpagebyaquarterinchanddrawscross-stylecropmarks.Itismostlyuse-fulwithtypesettersthatcansetthepagesizeautomatically. -KThisoptioncausescommentsinincludedPostScriptgraphics,fontfiles,andheaderstobe removed.ThisissometimesnecessarytogetaroundbugsinspoolersorPostScriptpostprocessingprograms.Specifically,the%%Pagecomments,whenleftin,oftencausedifficulties.Useofthisflagcancausesomeincludedgraphicstofail,sincethePostScript headermacrosfromsomesoftwarepackagesreadportionsoftheinputstreamlinebyline, searchingforaparticularcomment.Thisoptionhasbeenturnedoffbydefaultbecause PostScriptpreviewersandspoolershavebeengettingbetter. 1June1996 2 DVIPS(1) DVIPS(1) -lnum ThelastpageprintedwillbethefirstonenumberednumDefaultisthelastpageinthe document.Ifthenumisprefixedbyanequalssign,thenit(andanyargumenttothe-p option)istreatedasasequencenumber,ratherthanavaluetocomparewith\count0values.Thus,using-l=9willendwiththeninthpageofthedocument,nomatterwhatthe pagesareactuallynumbered. -mSpecifymanualfeedforprinter. -MTurnsofftheautomaticfontgenerationfacility.Ifanyfontsaremissing,commandsto generatethefontsareappendedtothefilemissfont.loginthecurrentdirectory;thisfile canthenbeexecutedanddeletedtocreatethemissingfonts. -nnum Atmostnumpageswillbeprinted.Defaultis100000. -NTurnsoffstructuredcomments;thismightbenecessaryonsomesystemsthattrytointer-pretPostScriptcommentsinweirdways,oronsomePostScriptprinters.Oldversionsof TranScriptinparticularcannothandlemodernEncapsulatedPostScript. -oname TheoutputwillbesenttofilenameIfnofilenameisgiven,thedefaultnameisfile.ps wherethe.dvifilewascalledfile.dvi;ifthisoptionisn’tgiven,anydefaultintheconfigurationfileisused.Ifthefirstcharacterofthesuppliedoutputfilenameisanexclamation mark,thentheremainderwillbeusedasanargumenttopopen;thus,specifying!lpras theoutputfilewillautomaticallyqueuethefileforprinting.Thisoptionalsodisablesthe automaticreadingofthePRINTERenvironmentvariable,andturnsofftheautomatic sendingofcontrolDifitwasturnedonwiththe-Foptionorintheconfigurationfile;use -Fafterthisoptionifyouwantboth. -Ooffset Movetheoriginbyacertainamount.Theoffsetisacomma-separatedpairofdimen-sions,suchas.1in,-.3cm(inthesamesyntaxusedinthepapersizespecial).Theoriginof thepageisshiftedfromthedefaultposition(ofoneinchdown,oneinchtotherightfrom theupperleftcornerofthepaper)bythisamount. -pnum Thefirstpageprintedwillbethefirstonenumberednum.Defaultisthefirstpageinthe document.Ifthenumisprefixedbyanequalssign,thenit(andanyargumenttothe-l option)istreatedasasequencenumber,ratherthanavaluetocomparewith\count0values.Thus,using-p=3willstartwiththethirdpageofthedocument,nomatterwhat thepagesareactuallynumbered. -pppagelist Acomma-separatedlistofpagesandranges(a-b)maybegiven,whichwillbeinterpreted as\count0values.Pagesnotspecifiedwillnotbeprinted.Multiple-ppoptionsmaybe specifiedorallpagesandpagerangescanbespecifiedwithone-ppoption. -Pprintername Setsuptheoutputfortheappropriateprinter.Thisisimplementedbyreadinginconfig.printername,whichcanthensettheoutputpipe(asin,!lpr-Pprinternameaswellas thefontpathsandanyotherconfig.psdefaultsforthatprinteronly.Notethatconfig.ps isreadbeforeconfig.printernameInaddition,anotherfilecalled˜ /.dvipsrcissearchedfor immediatelyafterconfig.ps;thisfileisintendedforuserdefaults.Ifno-Pcommandis given,theenvironmentvariablePRINTERischecked.Ifthatvariableexists,andacorrespondingconfigurationfileexists,thatconfigurationfileisreadin. -qRuninquietmode.Don’tchatteraboutpagesconverted,etc.;reportnothingbuterrors tostandarderror. 1June1996 3 DVIPS(1) DVIPS(1) -rStackpagesinreverseorder.Normally,page1willbeprintedfirst. -RRuninsecuremode.Thismeansthat‘‘backtick’’commandsfroma\special{}or\psf-file{}macrointhe(La)TeXsourcelike\special{psfile="‘zcatfoo.ps.Z"}or\psffile[7272 540720]{"‘zcatscreendump.ps.gz"}arenotexecuted. -sCausestheentireglobaloutputtobeenclosedinasave/restorepair.Thiscausesthefile tonotbetrulyconformant,andisthusnotrecommended,butisusefulifyouaredriving theprinterdirectlyanddon’tcaretoomuchabouttheportabilityoftheoutput. -Snum Setthemaximumnumberofpagesineach‘section’.Thisoptionismostcommonlyused withthe-ioption;seethatdocumentationaboveformoreinformation. -tpapertype Thissetsthepapertypetopapertype.Thepapertypeshouldbedefinedinoneoftheconfigurationfiles,alongwiththeappropriatecodetoselectit.(Currentlyknowntypes includeletter,legal,ledger,a4,a3,)Youcanalsospecify-tlandscape,whichrotatesa documentby90degrees.Torotateadocumentwhosesizeisnotletter,youcanusethe -toptiontwice,onceforthepagesize,andonceforlandscape.Theupperleftcornerof eachpageinthe.dvifileisplacedoneinchfromtheleftandoneinchfromthetop.Use ofthisoptionishighlydependentontheconfigurationfile.Notethatexecutingtheletter ora4orotherPostScriptoperatorscausethedocumenttobenonconformingandcan causeitnottoprintoncertainprinters,sothepapersizeshouldnotexecutesuchan operatorifatallpossible. -Toffset Setthepapersizetothegivenpairofdimensions.Thisoptiontakesitsargumentsinthe samestyleas-O.Itoverridesanypapersizespecialinthedvifile. -UDisableaPostScriptvirtualmemorysavingoptimizationthatstoresthecharactermetric informationinthesamestringthatisusedtostorethebitmapinformation.Thisisonly necessarywhendrivingtheXerox4045PostScriptinterpreter.Itiscausedbyabugin thatinterpreterthatresultsin‘garbage’onthebottomofeachcharacter.Notrecommendedunlessyoumustdrivethisprinter. -VDownloadnon-residentPostScriptfontsasbitmaps.Thisrequiresuseof‘gsftopk’or ‘pstopk’orsomeothersuchprogram(s)inordertogeneratetherequiredbitmapfonts; theseprogramsaresuppliedwithdvips. -xnum Setthemagnificationratiotonum/1000.Overridesthemagnificationspecifiedinthe.dvi file.Mustbebetween10and100000. -Xnum Setthehorizontalresolutionindotsperinchtonum. -Ynum Settheverticalresolutionindotsperinchtonum. -ZCausesbitmappedfontstobecompressedbeforetheyaredownloaded,therebyreducing thesizeofthePostScriptfont-downloadinginformation.Especiallyusefulathighresolu-tionsorwhenverylargefontsareused.Willslowdownprintingsomewhat,especiallyon early68000-basedPostScriptprinters. SEEALSO mf(1),afm2tfm(1),tex(1),latex(1),lpr(1),dvips.texi. ENVIRONMENT DvipskusesthesameenvironmentvariablesandalgorithmsforfindingfontfilesasTeXandits friendsdo.SeethedocumentationfortheKpathsealibraryfordetails.(Repeatingithereistoo cumbersome.) 1June1996 4 DVIPS(1) DVIPS(1) KPATHSEADEBUG:TraceKpathsealookups;setto-1forcompletetracing. PRINTER:seeabove. NOTES PostScriptisaregisteredtrademarkofAdobeSystemsIncorporated. AUTHOR TomasRokicki<rokicki@cs.stanford.edu>;extendedtovirtualfontsbyDonKnuth.Pathsearchingandconfigurationmodificationsbykb@mail.tug.org. 1June1996 5